please help me with the following question with a good explanation on how to work it out for me and my son, Thanks in advance! evaluate the expression: (1000!-999!)/998!
Answer by stanbon(75887)

evaluate the expression: (1000!-999!)/998!
------------------------------------
Note: 1000! = 1000*999!
and 999! = 999*998!
---------------------------
Rewrite the problem:
(1000*999! - 999!) / 998!
----------------------------
= (999!)(1000 - 1)/998!
= (999*998!)(999)/998!
Cancel the 998! common to the numerator the denominator.
= 999*999
= 998001
======================
Cheers,
Stan H.